### v2

# A Manifold Perspective on the Statistical Generalization of Graph Neural Networks

Abstract
Convolutional neural networks have been successfully extended to operate on graphs, giving rise to Graph Neural Networks (GNNs). GNNs combine information from adjacent nodes by successive applications of graph convolutions. GNNs have been implemented successfully in various learning tasks while the theoretical understanding of their generalization capability is still in progress. In this paper, we leverage manifold theory to analyze the statistical generalization gap of GNNs operating on graphs constructed on sampled points from manifolds. We study the generalization gaps of GNNs on both node-level and graph-level tasks. We show that the generalization gaps decrease with the number of nodes in the training graphs, which guarantees the generalization of GNNs to unseen points over manifolds. We validate our theoretical results in multiple real-world datasets.
1 Introduction
Graph convolutional neural networks (GNNs) [1, 2, 3] have emerged as one of the leading tools for processing graph-structured data. There is abundant evidence of their empirical success across various fields, including but not limited to weather prediction [4], protein structure prediction in biochemistry [5, 6], resource allocation in wireless communications [7], social network analysis in sociology [8], point cloud in 3D model reconstruction [9] and learning simulators [10].
The effectiveness of GNNs relies on their empirical ability to predict over unseen data. This capability is evaluated theoretically with statistical generalization in deep learning theory [11], which quantifies the difference between the empirical risk (i.e. training error) and the statistical risk (i.e. testing error). Despite the abundant evidence of GNNs’ generalization capabilities in practice, developing concrete theories to explain their generalization is an active area of research. Many recent works have studied the generalization bounds of GNNs without any dependence on the underlying model responsible for generating the graph data [12, 13, 14]. Generalization analysis on graph classification, when graphs are drawn from random limit models, is also studied in a series of works [15, 16, 17, 18]. In this work, we follow a similar methodology to formulate graph data on continuous topological spaces – specifically, manifolds – which are realistic models that enable rigorous theoretical analysis and a deeper understanding of the behaviors of GNNs.
Specifically, we explore the generalization gap of GNNs through the lens of manifold theory on both node level and graph level tasks. The graphs are constructed based on points randomly sampled from underlying manifolds, indicating that the manifold can be viewed as a statistical model for these discretely sampled points. As convolutional neural network architectures have been established over manifolds [19], the convergence of GNNs to manifold neural networks (MNNs) facilitates a detailed generalization analysis over GNNs. We demonstrate that, with an appropriate graph construction (Definition 2 or Definition 3), the generalization gap between empirical and statistical risks decreases with the number of sampled points in the graphs while increasing with the manifold dimension. The main findings of this paper (Theorem 1 and Theorem 2) provide a unified approach, showing that GNNs can effectively generalize to other unseen data over the manifold.
Theorem (Informal, Node classification on Gaussian kernel based graphs).
Consider a graph constructed with $N$ randomly sampled points over a $d$-dimensional manifold ${\mathcal{M}}$ with respect to the measure $\mu$ over the manifold. Then, with probability $1-\delta$ the generalization gap (GA) of a GNN with $L$ layers and $F$ features in each layer trained to predict the node labels on this graph satisfies
$GA\approx{\mathcal{O}}\left(\frac{LF^{L-1}}{\sqrt{N}}\left(\frac{\log 1/\delta}{N}\right)^{\frac{1}{d+4}}+F^{L}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}\right).$ (1)
Simulations on the ArXiv dataset [20] further validate our theoretical results as Figure 1 shows. Both the accuracy and loss generalization gaps decrease approximately linearly with the number of nodes in the logarithmic scale. We provide generalization bounds on both node and graph levels. For node level tasks, the generalization guarantees that a GNN trained on finitely sampled points over the manifold can accurately predict the labels of other unseen points over the manifold. For graph level tasks, the generalization assures that a GNN trained to classify graphs sampled from different manifolds can predict the labels of other unseen graphs sampled from these manifolds.
(a) Accuracy Generalization Gap
(b) Loss Generalization Gap
Figure 1: OBGN-Arxiv accuracy and loss generalization gaps. The GNN is trained over graphs with the number of nodes indicated on the x-axis, and the generalization gap (difference between the training and loss datasets) is measured, both in terms of the accuracy (Figure 1(a)), and the loss (Figure 1(b)) with which we trained – the cross-entropy loss. As can be seen, in both cases the generalization gap presents a logarithmically linear behavior with respect to the number of nodes.
2 Related works
2.1 Generalization bounds of GNNs
Node level tasks
We first give a brief recap of the generalization bounds of GNNs on node level tasks. In [12], the authors give a generalization bound of GNNs with a Vapnik–Chervonenkis dimension of GNNs. The authors in [14] analyze the generalization of a single-layer GNN based on stability analysis, which is further extended to a multi-layer GNN in [21]. In [22], the authors give a novel PAC-Bayesian analysis on the generalization bound of GNNs across arbitrary subgroups of training and testing datasets. The authors derive generalization bounds for GNNs via transductive uniform stability and transductive Rademacher complexity in [23, 24, 25]. The authors in [26] propose a size generalization analysis of GNNs correlated to the discrepancy between local distributions of graphs. Different from these works, we consider a continuous manifold model when generating the graph data, which is theoretically powerful and realistic when characterizing real-world data.
Graph level tasks
There are also related works on the generalization analysis of GNNs on graph level tasks. In [13], the authors form the generalization bound via Rademacher complexity. The authors in [27] build a PAC-Bayes framework to analyze the generalization capabilities of graph convolutional networks [28] and message-passing GNNs [29], based on which the authors in [30] improve the results and prove a lower bound. The works in [16, 17, 18] are most related to ours, which also consider the generalization of GNNs on a graph limit model. Different from our setting, the authors see the graph limit as a random continuous model. They study the generalization of graph classification problems with message-passing GNNs with graphs belonging to the same category sampled from a continuous limit model. The generalization bound grows with the model complexity and decreases with the number of nodes in the graph. We show that a GNN trained on a single graph sampled from each manifold is enough, and can generalize and classify unseen graphs sampled from the manifold set.
2.2 Neural networks on manifolds
Geometric deep learning has been proposed in [31] with neural network architectures raised in manifold space. The authors in [32] and [33] provide neural network architectures for manifold-valued data. In [34] and [19], the authors define convolutional operation over manifolds and see the manifold convolution as a generalization of graph convolution, which establishes the limit of neural networks on large-scale graphs as manifold neural networks (MNNs). The authors in [35] further establish the relationship between GNNs and MNNs with non-asymptotic convergence results for different graph constructions.
3 Preliminaries
3.1 Graph neural networks
Setup
An undirected graph ${\mathbf{G}}=({\mathcal{V}},{\mathcal{E}},{\mathcal{W}})$ contains a node set ${\mathcal{V}}$ with $N$ nodes and an edge set ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$. The weight function ${\mathcal{W}}:{\mathcal{E}}\rightarrow{\mathbb{R}}$ assigns weight values to the edges. We define the Graph Laplacian ${\mathbf{L}}=\text{diag}({\mathbf{A}}\textbf{1})-{\mathbf{A}}$ where ${\mathbf{A}}\in{\mathbb{R}}^{N\times N}$ is the adjacency matrix. Graph signals are defined as functions mapping nodes to a feature value ${\mathbf{x}}\in{\mathbb{R}}^{N}$.
Graph convolutions and frequency response
A graph convolutional filter ${\mathbf{h}}_{\mathbf{G}}$ is composed of consecutive graph shifts by graph Laplacian, defined as ${\mathbf{h}}_{\mathbf{G}}({\mathbf{L}}){\mathbf{x}}=\sum_{k=0}^{K-1}h_{k}{\mathbf{L}}^{k}{\mathbf{x}}$ with $\{h_{k}\}_{k=0}^{K-1}$ as filter parameters. We replace ${\mathbf{L}}$ with eigendecomposition ${\mathbf{L}}={\mathbf{V}}\bm{\Lambda}{\mathbf{V}}^{H}$, where ${\mathbf{V}}$ is the eigenvector matrix and $\bm{\Lambda}$ is a diagonal matrix with eigenvalues $\{\lambda_{i,N}\}_{i=1}^{N}$ as the entries. The spectral representation of a graph filter is
${\mathbf{V}}^{H}{\mathbf{h}}_{\mathbf{G}}({\mathbf{L}}){\mathbf{x}}=\sum_{k=1}^{K-1}h_{k}\bm{\Lambda}^{k}{\mathbf{V}}^{H}{\mathbf{x}}=\hat{h}(\bm{\Lambda}){\mathbf{V}}^{H}{\mathbf{x}}.$ (2)
This leads to a point-wise frequency response of the graph convolution as $\hat{h}(\lambda)=\sum_{k=0}^{K-1}h_{k}\lambda^{k}$.
Graph neural networks
A graph neural network (GNN) is a layered architecture, where each layer consists of a bank of graph convolutional filters followed by a point-wise nonlinearity $\sigma:{\mathbb{R}}\to{\mathbb{R}}$. Specifically, the $l$-th layer of a GNN that produces $F_{l}$ output features $\{{\mathbf{x}}_{l}^{p}\}_{p=1}^{F_{l}}$ with $F_{l-1}$ input features $\{{\mathbf{x}}^{q}_{l-1}\}_{q=1}^{F_{l-1}}$ is written as
${\mathbf{x}}_{l}^{p}=\sigma\left(\sum_{q=1}^{F_{l-1}}{\mathbf{h}}_{\mathbf{G}}^{lpq}({\mathbf{L}}){\mathbf{x}}^{q}_{l-1}\right),$ (3)
for each layer $l=1,2\cdots,L$. The graph filter ${\mathbf{h}}_{\mathbf{G}}^{lpq}({\mathbf{L}})$ maps the $q$-th feature of layer $l-1$ to the $p$-th feature of layer $l$. We denote the GNN as a mapping $\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}},{\mathbf{x}})$, where ${\mathbf{H}}\in{\mathcal{H}}\subset{\mathbb{R}}^{P}$ denotes a set of the graph filter coefficients at all layers and ${\mathcal{H}}$ denotes the set of all possible parameter sets.
3.2 Manifold neural networks
Setup
We consider a $d$-dimensional compact, smooth and differentiable Riemannian submanifold ${\mathcal{M}}$ embedded in ${\mathbb{R}}^{\mathsf{M}}$ with finite volume. This induces a Hausdorff probability measure $\mu$ over the manifold with density function $\rho:{\mathcal{M}}\rightarrow(0,\infty)$, assumed to be bounded as $0<\rho_{min}\leq\rho(x)\leq\rho_{max}<\infty$ for all $x\in{\mathcal{M}}$. The manifold data supported on each point $x\in{\mathcal{M}}$ is defined by scalar functions $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$ [34]. We use $L^{2}({\mathcal{M}})$ to denote $L^{2}$ functions over ${\mathcal{M}}$ with respect to measure $\mu$. The manifold with probability density function $\rho$ is equipped with a weighted Laplace operator [36], generalizing the Laplace-Beltrami operator as
${\mathcal{L}}_{\rho}f=-\frac{1}{2\rho}\text{div}(\rho^{2}\nabla f),$ (4)
with div denoting the divergence operator of ${\mathcal{M}}$ and $\nabla$ denoting the gradient operator of ${\mathcal{M}}$ [31, 37].
Manifold convolutions and frequency responses
The manifold convolution operation is defined relying on the Laplace operator ${\mathcal{L}}_{\rho}$ and on the heat diffusion process over the manifold [34]. For a function $f\in L^{2}({\mathcal{M}})$ as the initial heat condition over ${\mathcal{M}}$, the heat condition diffused by a unit time step can be explicitly written as $e^{-{\mathcal{L}}_{\rho}}f$. A manifold convolutional filter [34] can be defined in a diffuse-and-sum manner as
$g(x)={\mathbf{h}}({\mathcal{L}}_{\rho})f(x)=\sum_{k=0}^{K-1}h_{k}e^{-k{\mathcal{L}}_{\rho}}f(x)\text{,}$ (5)
with the $k$-th diffusion scaled with a filter parameter $h_{k}\in{\mathbb{R}}$. We consider the case in which the Laplace operator is self-adjoint with respect to the inner product defined in (24) and positive-semidefinite and the manifold ${\mathcal{M}}$ is compact, in this case, ${\mathcal{L}}_{\rho}$ has real, positive and discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$, written as ${\mathcal{L}}_{\rho}\bm{\phi}_{i}=\lambda_{i}\bm{\phi}_{i}$ where $\bm{\phi}_{i}$ is the eigenfunction associated with eigenvalue $\lambda_{i}$. The eigenvalues are ordered in increasing order as $0=\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots$, and the eigenfunctions are orthonormal and form an eigenbasis of $L^{2}({\mathcal{M}})$. When mapping a manifold signal onto the eigenbasis $[\hat{f}]_{i}=\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}$, the manifold convolution can be written in the spectral domain as
$[\hat{g}]_{i}=\sum_{k=0}^{K-1}h_{k}e^{-k\lambda_{i}}[\hat{f}]_{i}\text{.}$ (6)
Hence, the frequency response of manifold filter is given by $\hat{h}(\lambda)=\sum_{k=0}^{K-1}h_{k}e^{-k\lambda}$.
Manifold neural networks
A manifold neural network (MNN) is constructed by cascading $L$ layers, each of which contains a bank of manifold convolutional filters and a pointwise nonlinearity $\sigma:{\mathbb{R}}\rightarrow{\mathbb{R}}$. The output manifold function of each layer $l=1,2\cdots,L$ can be explicitly denoted as
$f_{l}^{p}(x)=\sigma\left(\sum_{q=1}^{F_{l-1}}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}(x)\right),$ (7)
where $f_{l-1}^{q}$, $1\leq q\leq F_{l-1}$ is the $q$-th input feature from layer $l-1$ and $f_{l}^{p}$, $1\leq p\leq F_{l}$ is the $p$-th output feature of layer $l$. We denote MNN as a mapping $\boldsymbol{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)$, where ${\mathbf{H}}\in{\mathcal{H}}\subset{\mathbb{R}}^{P}$ is a collective set of filter parameters in all the manifold convolutional filters.
4 Generalization analysis of GNNs based on manifold
We consider a manifold ${\mathcal{M}}$ as defined in Section 3.2, with a weighted Laplace operator ${\mathcal{L}}_{\rho}$ as (4). Since functions $f\in L^{2}({\mathcal{M}})$ characterize information over manifold ${\mathcal{M}}$, we restrict our analysis to a finite-dimensional subset of $L^{2}({\mathcal{M}})$ up to some eigenvalue of ${\mathcal{L}}_{\rho}$, defined as a bandlimited signal.
Definition 1.
A manifold signal $f\in L^{2}({\mathcal{M}})$ is bandlimited if there exists some $\lambda>0$ such that for all eigenpairs $\{\lambda_{i},\bm{\phi}_{i}\}_{i=1}^{\infty}$ of the weighted Laplacian ${\mathcal{L}}_{\rho}$ when $\lambda_{i}>\lambda$, we have $\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}=0$.
Suppose we are given a set of $N$ i.i.d. randomly sampled points $X_{N}=\{x_{i}\}_{i=1}^{N}$ over ${\mathcal{M}}$, with $x_{i}\in{\mathcal{M}}$ sampled according to measure $\mu$. We construct a graph ${\mathbf{G}}({\mathcal{V}},{\mathcal{E}},{\mathcal{W}})$ on these $N$ sampled points $X_{N}$, where each point $x_{i}$ is a vertex of graph ${\mathbf{G}}$, i.e. ${\mathcal{V}}=X_{N}$. Each pair of vertices $(x_{i},x_{j})$ is connected with an edge while the weight attached to the edge ${\mathcal{W}}(x_{i},x_{j})$ is determined by a kernel function $K_{\epsilon}$. The kernel function is decided by the Euclidean distance $\|x_{i}-x_{j}\|$ between these two points. The graph Laplacian denoted as ${\mathbf{L}}_{N}$ can be calculated based on the weight function [38]. The constructed graph Laplacian with an appropriate kernel function has been proved to approximate the Laplace operator ${\mathcal{L}}_{\rho}$ of ${\mathcal{M}}$ [39, 40, 41]. We present the following two definitions of $K_{\epsilon}$.
Definition 2 (Gaussian kernel based graph [40]).
The graph ${\mathbf{G}}(X_{N},{\mathcal{E}},{\mathcal{W}})$ can be constructed as a dense graph degree when the kernel function is defined as
${\mathcal{W}}(x_{i},x_{j})=K_{\epsilon}\left(\frac{\|x_{i}-x_{j}\|^{2}}{\epsilon}\right)=\frac{1}{N}\frac{1}{\epsilon^{d/2+1}(4\pi)^{d/2}}e^{-\frac{\|x_{i}-x_{j}\|^{2}}{4\epsilon}},\quad(x_{i},x_{j})\in{\mathcal{E}}.$ (8)
The weight function of a Gaussian kernel based graph is defined on unbounded support (i.e. $[0,\infty)$), which connects $x_{i}$ and $x_{j}$ regardless of the distance between them. This results in a dense graph with $N^{2}$ edges. In particular, this Gaussian kernel based graph has been widely used to define the weight value function due to the good approximation properties of the corresponding graph Laplacians to the manifold Laplace operator [41, 40, 42].
Definition 3 ($\epsilon$-graph [39]).
The graph ${\mathbf{G}}(X_{N},{\mathcal{E}},{\mathcal{W}})$ can be constructed as an $\epsilon$-graph with the kernel function defined as
${\mathcal{W}}(x_{i},x_{j})=K_{\epsilon}\left(\frac{\|x_{i}-x_{j}\|^{2}}{\epsilon}\right)=\frac{1}{N}\frac{d+2}{\epsilon^{d/2+1}\alpha_{d}}\mathbbm{1}_{[0,1]}\left(\frac{\|x_{i}-x_{j}\|^{2}}{\epsilon}\right),\quad(x_{i},x_{j})\in{\mathcal{E}},$ (9)
where $\alpha_{d}$ is the volume of a unit ball of dimension $d$ and $\mathbbm{1}$ is the characteristic function.
The weight function of an $\epsilon$-graph is defined on a bounded support, i.e., only nodes that are within a certain distance of one another can be connected by an edge. It has also been shown to provide a good approximation of the manifold Laplace operator [39]. Figure 2 gives an illustration of both Gaussian kernel based graphs and $\epsilon$-graphs sampled from point cloud models [43].
(a) Gaussian kernel based graphs
(b) $\epsilon$-graphs
Figure 2: Illustration of the constructed graphs on points sampled over a chair and a table model.
4.1 Manifold label prediction via node label prediction
Suppose we have an input manifold signal $f\in L^{2}({\mathcal{M}})$ and a label (i.e. target) manifold signal $g\in L^{2}({\mathcal{M}})$ over ${\mathcal{M}}$. With an MNN $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},\cdot)$, we predict the target value $g(x)$ based on input $f(x)$ at each point $x\in{\mathcal{M}}$. By sampling $N$ points $X_{N}$ over this manifold, we can approximate this problem in a discrete graph domain. Consider a graph ${\mathbf{G}}(X_{N},{\mathcal{E}},{\mathcal{W}})$ constructed with $X_{N}$ as either a Gaussian kernel based graph (Definition 2) or an $\epsilon$-graph (Definition 3) equipped with the graph Laplacian ${\mathbf{L}}_{N}$. Suppose we are given graph signal $\{{\mathbf{x}},{\mathbf{y}}\}$ sampled from $\{f,g\}$ to train a GNN $\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},\cdot)$, explicitly written as
$[{\mathbf{x}}]_{i}=f(x_{i}),\qquad[{\mathbf{y}}]_{i}=g(x_{i})\quad\text{for all }x_{i}\in X_{N}.$ (10)
We assume that the filters in MNN $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},\cdot)$ and GNN $\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},\cdot)$ are low-pass filters, which are defined explicitly as follows.
Definition 4.
A filter is a low-pass filter if its frequency response satisfies
$\left|\hat{h}(\lambda)\right|={\mathcal{O}}\left(\lambda^{-d}\right),\quad\left|\hat{h}^{\prime}(\lambda)\right|\leq C_{L}\lambda^{-d-1},\quad\lambda\in(0,\infty).$ (11)
To introduce the first of our two main results, we require the introduction of two assumptions.
AS 1.
(Normalized Lipschitz nonlinearity) The nonlinearity $\sigma$ is normalized Lipschitz continuous, i.e., $|\sigma(a)-\sigma(b)|\leq|a-b|$, with $\sigma(0)=0$.
AS 2.
(Normalized Lipschitz loss function) The loss function $\ell$ is normalized Lipschitz continuous, i.e., $|\ell(y_{i},y)-\ell(y_{j},y)|\leq|y_{i}-y_{j}|$, with $\ell(y,y)=0$.
Assumption 1 is satisfied by most activations used in practice such as ReLU, modulus and sigmoid.
The generalization gap is evaluated between the empirical risk over the discrete graph model and the statistical risk over manifold model, with the manifold model viewed as a statistical model since the expectation of the sampled point is with respect to the measure $\mu$ over the manifold. The empirical risk over the sampled graph that we trained to minimize is therefore defined as
$R_{\mathbf{G}}({\mathbf{H}})=\frac{1}{N}\sum_{i=1}^{N}\ell\left([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i}\right).$ (12)
The statistical risk over the manifold is defined as
$R_{\mathcal{M}}({\mathbf{H}})=\int_{\mathcal{M}}\ell\left(\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),g(x)\right)\text{d}\mu(x).$ (13)
The generalization gap is defined to be
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}\left|R_{\mathcal{M}}({\mathbf{H}})-R_{\mathbf{G}}({\mathbf{H}})\right|.$ (14)
Theorem 1.
Suppose the GNN and MNN with low-pass filters (Definition 4) have $L$ layers with $F$ features in each layer and the input signal is bandlimited (Definition 1). Under Assumptions 1 and 2 it holds in probability at least $1-\delta$ that
$GA\leq LF^{L-1}\left(C_{1}\frac{\epsilon}{\sqrt{N}}+C_{2}\frac{\sqrt{\log(1/\delta)}}{N}\right)+F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}$ (15)
  1. 1.
    When the graph is constructed with a Gaussian kernel (8), then $\epsilon\sim\left(\frac{\log(C/\delta)}{N}\right)^{\frac{1}{d+4}}$.
  2. 2.
    When the graph is constructed as an $\epsilon$-graph as (9), then $\epsilon\sim\left(\frac{\log(CN/\delta)}{N}\right)^{\frac{1}{d+4}}$.
Proof.
See Appendix D for proof and the definitions of $C_{1}$, $C_{2}$ and $C_{3}$. ∎
Theorem 1 shows that the generalization gap decreases approximately linearly with the number of nodes $N$ in the logarithmic scale. It also increases with the dimension of the underlying manifold $d$. Another observation is that the generalization gap scales with the size of GNN architectures. Specifically, it scales polynomially with the number of features $F$, while the exponent depends on the number of layers $L$. This generalization result attests that when the graph is supported on a manifold, a GNN trained on this discrete graph model can approximate the MNN performance. GNNs trained on these finite sampled points can generalize to predict the labels of other unseen points over the manifold with large enough nodes in the graph. To generalize effectively over a manifold with a higher dimension, we need more sampled points to construct the graph to guarantee the generalization.
4.2 Manifold classification via graph classification
Suppose we have a set of manifolds $\{{\mathcal{M}}_{k}\}_{k=1}^{K}$, each of which is $d_{k}$-dimensional, smooth, compact, differentiable and embedded in ${\mathbb{R}}^{\mathsf{M}}$ with measure $\mu_{k}$. Each manifold ${\mathcal{M}}_{k}$ equipped with a weighted Laplace operator ${\mathcal{L}}_{\rho,k}$ is labeled with $y_{k}\in{\mathbb{R}}$. We assume to have access to $N_{k}$ randomly sampled points according to measure $\mu_{k}$ over each manifold ${\mathcal{M}}_{k}$ and construct $K$ graphs $\{{\mathbf{G}}_{k}\}_{k=1}^{K}$ with graph Laplacian ${\mathbf{L}}_{N,k}$. The GNN $\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,\cdot},{\mathbf{x}}_{\cdot})$ is trained on this set of graphs with ${\mathbf{x}}_{k}$ as the input graph signal sampled from the manifold signal $f_{k}\in L^{2}({\mathcal{M}}_{k})$ and $y_{k}\in{\mathbb{R}}$ as the scalar target label. The final output of the GNN is set to be the average of the output signal values on each node while the output of MNN $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,\cdot},f_{\cdot})$ is the statistical average value of the output signal over the manifold. A loss function $\ell$ evaluates the difference between the output of GNN and MNN with the target label. The empirical risk of the GNN is
$R_{\mathbf{G}}({\mathbf{H}})=\sum_{k=1}^{K}\ell\left(\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}[\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right).$ (16)
While the output of MNN is the average value over the manifold, the statistical risk is defined based on the loss evaluated between the MNN output and the label as
$R_{\mathcal{M}}({\mathbf{H}})=\sum_{k=1}^{K}\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x)\text{d}\mu_{k}(x),y_{k}\right).$ (17)
The generalization gap is therefore
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}\left|R_{\mathcal{M}}({\mathbf{H}})-R_{\mathbf{G}}({\mathbf{H}})\right|.$ (18)
Theorem 2.
Suppose the GNN and MNN with low-pass filters (Definition 4) have $L$ layers with $F$ features in each layer and the input signal is bandlimited (Definition 1). Under Assumptions 1 and 2 it holds in probability at least $1-\delta$ that
$GA\leq LF^{L-1}\sum_{k=1}^{K}\left(\frac{C_{1}}{\sqrt{N_{k}}}\epsilon+C_{2}\frac{\sqrt{\log(1/\delta)}}{N_{k}}\right)+F^{L}C_{3}\sum_{k=1}^{K}\left(\frac{\log N_{k}}{N_{k}}\right)^{\frac{1}{d_{k}}}$ (19)
  1. 1.
    When the graph is constructed with a Gaussian kernel (8), then $\epsilon\sim\left(\frac{\log(C/\delta)}{N_{k}}\right)^{\frac{1}{d_{k}+4}}$.
  2. 2.
    When the graph is constructed as an $\epsilon$-graph as (9), then $\epsilon\sim\left(\frac{\log(CN_{k}/\delta)}{N_{k}}\right)^{\frac{1}{d_{k}+4}}$.
Proof.
See Appendix E for proof and the definitions of $C_{1}$, $C_{2}$ and $C_{3}$. ∎
Theorem 2 shows that a single graph sampled from the underlying manifold with large enough sampled points $N$ can provide an effective approximation to classify the manifold itself. The generalization gap also attests that the trained GNN can generalize to classify other unseen graphs sampled from the same manifold. Similar to the generalization result in node-level tasks, the generalization gap decreases with the number of points sampled over each manifold while increasing with the manifold dimension. A higher dimensional manifold, i.e. higher complexity, needs more samples to guarantee the generalization. The generalization gap also scales with the size of GNN architecture polynomially with the number of features and exponentially with the number of layers.
(a) Two Layers
(b) Three Layers
(c) Two Layers
(d) Three Layers
Figure 3: In the top row, we plot the difference in the accuracy and loss for columns 3(a), 3(b) and 3(c), 3(d), respectively. On the bottom row, we plot the actual train and test values of the accuracy (3(a), 3(b)), and the loss (3(c), 3(d)). The plots are for the Arxiv dataset and $\{64,128,256,512\}$ hidden units. For the bottom row, we also calculate the linear fit for the values whose training accuracy is below $95\%$, showing that our linear bound on the logarithm of the generalization gap for the logarithm of the number of nodes shares the same rate shown in Theorem 1.
5 Experiments
Dataset
Pearson
Correlation
Coefficient
OGBN-Arxiv $-0.9980$
Cora $-0.9686$
CiteSeer $-0.9534$
PubMed $-0.9761$
CS $-0.8969$
Physics $-0.9145$
Amazon $-0.9972$
Roman $-1.0000$
Table 1: Pearson correlation for $2$ layer, $64$ hidden units GNNs measured on the accuracy generalization gap.
Node classification
[…]
(a) Cora
(b) CiteSeer
(c) PubMed
Figure 4: Generalization gap as a function of the number of nodes in the training set for accuracy (top) and loss (bottom) for the Cora, CiteSeer, and PubMed datasets.
Our main goal is to show that the rate presented in Theorem 1 holds in practice. That is to say, if we plot the logarithm of the generalization gap as a function of the logarithm of the number of nodes we see a linear rate. To be consistent with the theory, we also want to show that if the number of layers or the size of the features increases, so does the generalization gap.
In Figure 3, we plot the generalization gap of the accuracy in the logarithmic scale for a two-layered GNN (Figure 3(a)), and for a three-layered GNN (Figure 3(b)). On the upper side, we can see that the generalization bound decreases with the number of nodes and that outside of the strictly overfitting regime (when the training loss is below $95\%$), the generalization gap shows a linear decay, as depicted in the dashed line. The same behavior can be seen in plots 3(c), and 3(d) which correspond to the loss for $2$ and $3$ layered GNNs. As predicted by our theory, the generalization gap increases with the number of features and layers in the GNN (see also Figure 1). The behavior of the training and testing accuracy as a function of the number of nodes is intuitive. For the training loss, when the number of nodes in the training set is small, the GNN can overfit the training data. As the number of features increases, the GNN’s capacity to overfit also increases.
(a) Amazon-Ratings
(b) Roman-Empire
(c) CoAuthors CS
(d) CoAuthors Physics
Figure 5: Accuracy generalization gap as a function of the number of nodes in the training set for the Amazon-Ratings, Roman-Empire, CoAuthors CS, and CoAuthors Physics datasets.
In Figures 4, and 5 we present the accuracy generalization gaps for $2$ and $3$ layers with $32$ and $64$ features. In the overfitting regime, the rate of our generalization bound seems to hold – decreases linearly with the number of nodes in the logarithmic scale. In the non-overfitting regime, our rate holds for the points whose training accuracy is below $95\%$. Also, we validate that the bound increases both with the number of features, as well as the number of layers.
In Table 1 we present the Pearson correlation coefficient to measure the linear relationships in the generalization gaps of a $2$ layers GNN with $64$ hidden units in all datasets considered. In almost every case, the coefficient is above $0.95$ which translates into a strong linear correlation. In Appendix F we explain how we computed these values. As seen in the experiment, the GNN generalization gap experiences a linear decay with respect to the number of nodes in the logarithmic scale. The bound presented in Theorem 1 is an upper bound on the generalization gap, whose rate can be seen to match the one seen in practice both for the loss, as well as the accuracy gaps.
Graph classification
(a) Differences of the outputs of 3-layer GNNs.
(b) Differences of the outputs of 4-layer GNNs.
Figure 6: Generalization gap as a function of number of nodes for average GNN output differences for graph classification over ModelNet10.
We evaluate the generalization gap on graph prediction using the ModelNet10 dataset [43]. We set the coordinates of each point as input graph signals, and the weights of the edges are calculated based on the Euclidean distance between the nodes. The generalization gap is calculated by training GNNs on graphs with $N=10,20,\dots,80$ sampled points, and plotting the differences between the average output of the trained GNNs on the trained graph and a testing graph with size $N=100$. Figure 6(a) shows the generalization gaps for GNNs with 3 layers and Figure 6(b) shows the results of GNNs with 4 layers. We can see that the output differences between the GNNs decrease with the number of nodes and scale with the size of the GNN architectures. This verifies the claims of Theorem 2. In Appendix F, we present experiment results on more model datasets.
6 Conclusion
We study the statistical generalization of GNNs from a manifold perspective. We considered graphs sampled from manifolds and proved that GNNs can effectively generalize to unseen data from the manifolds when the number of sampled points is large enough. We validate our results both on node-level and graph-level classification tasks with real-world datasets. The impact of this paper is to show a better understanding of GNN generalization capabilities from a statistical perspective relying on a continuous model. Our work also motivates the practical design of large-scale GNNs. For future work, we will study the switching behavior between the over-fitting, and non-over-fitting regimes of node classification.
Appendix A Induced manifold signals
The graph signal attached to this constructed graph ${\mathbf{G}}$ can be seen as the discretization of the continuous function over the manifold. Suppose $f\in L^{2}({\mathcal{M}})$, the graph signal ${\mathbf{x}}_{N}$ is composed of discrete data values of the function $f$ evaluated at $X_{N}$, i.e. $[{\mathbf{x}}_{N}]_{i}=f(x_{i})$ for $i=1,2\cdots,N$. With a sampling operator ${\mathbf{P}}_{N}:L^{2}({\mathcal{M}})\rightarrow L^{2}(X_{N})$, the discretization can be written as
${\mathbf{x}}_{N}={\mathbf{P}}_{N}f.$ (20)
Let $\mu_{N}$ be the empirical measure of the random sample as
$\mu_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}.$ (21)
Let $\{V_{i}\}_{i=1}^{N}$ be the decomposition [48] of ${\mathcal{M}}$ with respect to $X_{N}$ with $V_{i}\subset B_{r}(x_{i})$, where $B_{r}(x_{i})$ denoted the closed metric ball of radius $r$ centered at $x_{i}\in{\mathcal{M}}$. The decomposition can be achieved by the optimal transportation map $T:{\mathcal{M}}\rightarrow X_{N}$, which is defined by the $\infty$-Optimal Transport distance between $\mu$ and $\mu_{N}$.
$d_{\infty}(\mu,\mu_{N}):=\min_{T:T_{\#}\mu=\mu_{N}}\text{esssup}_{x\in{\mathcal{M}}}d(x,T(x)),$ (22)
where $T_{\#}\mu=\mu_{N}$ indicates that $\mu(T^{-1}(V))=\mu_{N}(V)$ for every $V_{i}$ of ${\mathcal{M}}$. This transportation map $T$ induces the partition $V_{1},V_{2},\cdots V_{N}$ of ${\mathcal{M}}$, where $V_{i}:=T^{-1}(\{x_{i}\})$ with $\mu(V_{i})=\frac{1}{N}$ for all $i=1,\cdots N$. The radius of $V_{i}$ can be bounded as $r\leq A(\log N/N)^{1/d}$ with $A$ related to the geometry of ${\mathcal{M}}$ [48, Theorem 2].
The manifold function induced by the graph signal ${\mathbf{x}}_{N}$ over the sampled graph ${\mathbf{G}}$ is defined by
$({\mathbf{I}}_{N}{\mathbf{x}}_{N})(x)=\sum_{i=1}^{N}[{\mathbf{x}}]_{i}\mathbbm{1}_{x\in V_{i}},\text{for all}\;x\in{\mathcal{M}}$ (23)
where we denote ${\mathbf{I}}_{N}:L^{2}(X_{N})\rightarrow L^{2}({\mathcal{M}})$ as the inducing operator.
Appendix B Convergence of GNN to MNN
The convergence of GNN on sampled graphs to MNN provides the support for the generalization analysis. We first introduce the inner product over the manifold. The inner product of signals $f,g\in L^{2}({\mathcal{M}})$ is defined as
$\langle f,g\rangle_{{\mathcal{M}}}=\int_{\mathcal{M}}f(x)g(x)\text{d}\mu(x),$ (24)
where $\text{d}\mu(x)$ is the volume element with respect to the measure $\mu$ over ${\mathcal{M}}$. Similarly, the norm of the manifold signal $f$ is
$\|f\|^{2}_{{\mathcal{M}}}={\langle f,f\rangle_{{\mathcal{M}}}}.$ (25)
Proposition 1.
Let ${\mathcal{M}}\subset{\mathbb{R}}^{\mathsf{M}}$ be an embedded manifold with weighted Laplace operator ${\mathcal{L}}_{\rho}$ and a bandlimited manifold signal $f$. Graph ${\mathbf{G}}_{N}$ is constructed based on a set of $N$ i.i.d. randomly sampled points $X_{N}=\{x_{1},x_{2},\cdots,x_{N}\}$ according to measure $\mu$ over ${\mathcal{M}}$. A graph signal ${\mathbf{x}}$ is the sampled manifold function values at $X_{N}$. The graph Laplacian ${\mathbf{L}}_{N}$ is calculated based on (8) or (9) with $\epsilon$ as the graph parameter. Let $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},\cdot)$ be a MNN on ${\mathcal{M}}$ (7) with $L$ layers and $F$ features in each layer. Let $\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},\cdot)$ be the GNN with the same architecture applied to the graph ${\mathbf{G}}_{N}$. Then, with the filters as low-pass and nonlinearities as normalized Lipschitz continuous, it holds in probability at least $1-\delta$ that
$\frac{1}{N}\sum_{i=1}^{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{2}\leq LF^{L-1}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (26)
where $C_{1}$, $C_{2}$ are constants defined in the following proof.
Proposition 2.
[35, Proposition 2, Proposition 4] Let ${\mathcal{M}}\subset{\mathbb{R}}^{\mathsf{M}}$ be equipped with Laplace operator ${\mathcal{L}}_{\rho}$, whose eigendecomposition is given by $\{\lambda_{i},\bm{\phi}_{i}\}_{i=1}^{\infty}$. Let ${\mathbf{L}}_{N}$ be the discrete graph Laplacian of graph weights defined as (8) (or (9)), with spectrum $\{\lambda_{i,N},\bm{\phi}_{i,N}\}_{i=1}^{N}$. Fix $K\in\mathbb{N}^{+}$ and assume that $\epsilon=\epsilon(N)\geq\left({\log(C/\delta)}/{N}\right)^{1/(d+4)}$ (or $\epsilon=\epsilon(N)\geq\left({\log(CN/\delta)}/{N}\right)^{1/(d+4)}$). Then, with probability at least $1-\delta$, we have
$|\lambda_{i}-\lambda_{i,N}|\leq C_{{\mathcal{M}},1}\lambda_{i}{\epsilon},\quad\|a_{i}\bm{\phi}_{i,N}-\bm{\phi}_{i}\|\leq C_{{\mathcal{M}},2}\frac{\lambda_{i}}{\theta_{i}}{\epsilon},$ (27)
with $a_{i}\in\{-1,1\}$ for all $i<K$ and $\theta$ the eigengap of ${\mathcal{L}}$, i.e., $\theta_{i}=\min\{\lambda_{i}-\lambda_{i-1},\lambda_{i+1}-\lambda_{i}\}$. The constants $C_{{\mathcal{M}},1}$, $C_{{\mathcal{M}},2}$ depend on $d$ and the volume, the injectivity radius and sectional curvature of ${\mathcal{M}}$.
Proof.
Because $\{x_{1},x_{2},\cdots,x_{N}\}$ is a set of randomly sampled points from ${\mathcal{M}}$, based on Theorem 19 in [49] we can claim that
$\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|=O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right).$ (28)
This also indicates that
$\left|\|{\mathbf{P}}_{N}f\|^{2}-\|f\|^{2}_{\mathcal{M}}\right|=O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right),$ (29)
which indicates $\|{\mathbf{P}}_{N}f\|=\|f\|_{\mathcal{M}}+O((\log(1/\delta)/N)^{1/4})$. We suppose that the input manifold signal is $\lambda_{M}$-bandlimited with $M$ spectral components. We first write out the filter representation as
$\|{\mathbf{h}}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f-{\mathbf{P}}_{N}{\mathbf{h}}({\mathcal{L}}_{\rho})f\|=\Bigg{\|}\sum_{i=1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\Bigg{\|}$ (30)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}+\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\Bigg{\|}$ (31)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\Bigg{\|}+\left\|\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\right\|$ (32)
The first part of (32) can be decomposed with the triangle inequality as
$\left\|\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\right\|$
$\leq\left\|\sum_{i=1}^{M}\left(\hat{h}(\lambda_{i,N})-\hat{h}(\lambda_{i})\right)\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\right\|+\left\|\sum_{i=1}^{M}\hat{h}(\lambda_{i})\left(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\right)\right\|.$ (33)
In equation (33), the first part relies on the difference of eigenvalues and the second part depends on the eigenvector difference. The first term in (33) is bounded with Cauchy-Schwartz inequality as
$\Bigg{\|}\sum_{i=1}^{M}(\hat{h}(\lambda_{i,n})-\hat{h}(\lambda_{i}))\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle$ $\bm{\phi}_{i,N}\Bigg{\|}\leq\sum_{i=1}^{M}\left|\hat{h}(\lambda_{i,N})-\hat{h}(\lambda_{i})\right||\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle|$ (34)
$\leq\|{\mathbf{P}}_{N}f\|\sum_{i=1}^{M}|\hat{h}^{\prime}(\lambda_{i})||\lambda_{i,N}-\lambda_{i}|$ (35)
$\leq\|{\mathbf{P}}_{N}f\|\sum_{i=1}^{M}C_{{\mathcal{M}},1}C_{L}\epsilon\lambda_{i}^{-d}$ (36)
$\leq\|{\mathbf{P}}_{N}f\|C_{L}C_{{\mathcal{M}},1}\epsilon\sum_{i=1}^{M}i^{-2}$ (37)
$\leq\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)C_{{\mathcal{M}},1}\epsilon\frac{\pi^{2}}{6}:=A_{1}(N)$ (38)
In (36), it depends on the low-pass filter assumption in Definition 4. In (37), we implement Weyl’s law [50] which indicates that eigenvalues of Laplace operator scales with the order $\lambda_{i}\sim i^{2/d}$. The last inequality comes from the fact that $\sum_{i=1}^{\infty}i^{-2}=\frac{\pi^{2}}{6}$. The second term in (33) can be bounded with the triangle inequality as
$\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i})\left(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\right)\Bigg{\|}$
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i})\left(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle{\mathbf{P}}_{N}\bm{\phi}_{i}\right)\Bigg{\|}$
$\qquad\qquad\qquad+\left\|\sum_{i=1}^{M}\hat{h}(\lambda_{i})\left(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle{\mathbf{P}}_{N}\bm{\phi}_{i}-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}{\mathbf{P}}_{N}\bm{\phi}_{i}\right)\right\|$ (39)
The first term in (39) can be bounded with inserting the eigenfunction convergence result in Proposition 2 as
$\left\|\sum_{i=1}^{M}\hat{h}(\lambda_{i})\left(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\right)\right\|$
$\leq\sum_{i=1}^{M}\left|\hat{h}(\lambda_{i})\right|\|{\mathbf{P}}_{N}f\|\|\bm{\phi}_{i,N}-{\mathbf{P}}_{N}\bm{\phi}_{i}\|$ (40)
$\leq\sum_{i=1}^{M}(\lambda_{i}^{-d})\frac{C_{{\mathcal{M}},2}\epsilon}{\theta_{i}}\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (41)
$\leq C_{{\mathcal{M}},2}\epsilon\frac{\pi^{2}}{6}\max_{i=1,\cdots,M}{\theta_{i}^{-1}}\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (42)
$:=A_{2}(M,N).$ (43)
Considering the low-pass filter assumption, the second term in (39) can be written as
$\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle{\mathbf{P}}_{N}\bm{\phi}_{i}-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}{\mathbf{P}}_{N}\bm{\phi}_{i})\Bigg{\|}$
$\leq\sum_{i=1}^{M}\left|\hat{h}(\lambda_{i,N})\right|\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\|{\mathbf{P}}_{N}\bm{\phi}_{i}\|$ (44)
$\leq\sum_{i=1}^{M}(\lambda_{i,N}^{-d})\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (45)
$\leq\sum_{i=1}^{M}(1+C_{{\mathcal{M}},1}\epsilon)^{-d}(\lambda_{i}^{-d})\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (46)
$\leq\frac{\pi^{2}}{6}\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right):=A_{3}(N)$ (47)
The term $\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ can be decomposed by inserting a term $\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle$ as
$\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ $\leq\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle+\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (48)
$\leq\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle\right|+\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (49)
$\leq\|{\mathbf{P}}_{N}f\|\|\bm{\phi}_{i,N}-{\mathbf{P}}_{N}\bm{\phi}_{i}\|+\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (50)
$\leq\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)\frac{C_{{\mathcal{M}},2}\lambda_{i}\epsilon}{\theta_{i}}+\sqrt{\frac{\log(1/\delta)}{N}}$ (51)
Then equation (46) can be bounded as
$\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})(\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle{\mathbf{P}}_{N}\bm{\phi}_{i}-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}{\mathbf{P}}_{N}\bm{\phi}_{i})\Bigg{\|}$
$\leq\sum_{i=1}^{M}(1+C_{{\mathcal{M}},1}\epsilon)^{-d}(\lambda_{i}^{-d})\left(\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)\frac{C_{{\mathcal{M}},2}\lambda_{i}\epsilon}{\theta_{i}}+\sqrt{\frac{\log(1/\delta)}{N}}\right)\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (52)
$\leq\frac{\pi^{2}}{6}\max_{i=1,\cdots,M}\frac{C_{{\mathcal{M}},2}\epsilon}{\theta_{i}}\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)+\frac{\pi^{2}}{6}\sqrt{\frac{\log(1/\delta)}{N}}$ (53)
The second term in (32) can be bounded with the eigenvalue difference bound in Proposition 2 as
$\left\|\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\right\|$ $\leq\sum_{i=M+1}^{N}(\lambda_{i,N}^{-d})\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (54)
$\leq\sum_{i=M+1}^{\infty}(\lambda_{i,N}^{-d})\|f\|_{\mathcal{M}}$ (55)
$\leq(1+C_{{\mathcal{M}},1}\epsilon)^{-d}\sum_{i=M+1}^{\infty}(\lambda_{i}^{-d})\|f\|_{\mathcal{M}}$ (56)
$\leq M^{-1}\|f\|_{\mathcal{M}}:=A_{4}(M).$ (57)
We note that the bound is made up by terms $A_{1}(N)+A_{2}(M,N)+A_{3}(N)+A_{4}(M)$, related to the bandwidth of manifold signal $M$ and the number of sampled points $N$. This makes the bound scale with the order
$\|{\mathbf{h}}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f-{\mathbf{P}}_{N}{\mathbf{h}}({\mathcal{L}}_{\rho})f\|\leq C_{1}^{\prime}\epsilon+C_{2}^{\prime}\epsilon\theta_{M}^{-1}+C_{3}^{\prime}\sqrt{\frac{\log(1/\delta)}{N}}+C_{4}^{\prime}M^{-1},$ (58)
with $C_{1}^{\prime}=C_{L}C_{{\mathcal{M}},1}\frac{\pi^{2}}{6}\|f\|_{\mathcal{M}}$, $C_{2}^{\prime}=C_{{\mathcal{M}},2}\frac{\pi^{2}}{6}$, $C_{3}^{\prime}=\frac{\pi^{2}}{6}$ and $C_{4}^{\prime}=\|f\|_{\mathcal{M}}$. As $N$ goes to infinity, for every $\delta>0$, there exists some $M_{0}$, such that for all $M>M_{0}$ it holds that $A_{4}(M)\leq\delta/2$. There also exists $n_{0}$, such that for all $N>n_{0}$, it holds that $A_{1}(N)+A_{2}(M_{0},N)+A_{3}(N)\leq\delta/2$. We can conclude that the summations converge as $N$ goes to infinity. We see $M$ large enough to have $M^{-1}\leq\delta^{\prime}$, which makes the eigengap $\theta_{M}$ also bounded by some constant. We combine the first two terms as
$\|{\mathbf{h}}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f-{\mathbf{P}}_{N}{\mathbf{h}}({\mathcal{L}}_{\rho})f\|\leq C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}},$ (59)
with $C_{1}=C_{1}^{\prime}+C_{2}^{\prime}\theta^{-1}_{\delta^{\prime-1}}$ and $C_{2}=C_{3}^{\prime}$ To bound the output difference of MNNs, we need to write in the form of features of the final layer
$\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{P}}_{N}f)-{\mathbf{P}}_{N}\bm{\Phi}$ $({\mathbf{H}},{\mathcal{L}}_{\rho},f))\|=\left\|\sum_{q=1}^{F}{\mathbf{x}}_{n,L}^{q}-\sum_{q=1}^{F}{\mathbf{P}}_{N}f_{L}^{q}\right\|\leq\sum_{q=1}^{F}\left\|{\mathbf{x}}_{n,L}^{q}-{\mathbf{P}}_{N}f_{L}^{q}\right\|.$ (60)
By inserting the definitions, we have
$\left\|{\mathbf{x}}_{n,l}^{p}-{\mathbf{P}}_{N}f_{l}^{p}\right\|=\left\|\sigma\left(\sum_{q=1}^{F}{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}\right)-{\mathbf{P}}_{N}\sigma\left(\sum_{q=1}^{F}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\right)\right\|$ (61)
with ${\mathbf{x}}_{n,0}={\mathbf{P}}_{N}f$ as the input of the first layer. With a normalized point-wise Lipschitz nonlinearity, we have
$\|{\mathbf{x}}_{n,l}^{p}-{\mathbf{P}}_{N}f_{l}^{p}$ $\|\leq\left\|\sum_{q=1}^{F}{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}-{\mathbf{P}}_{N}\sum_{q=1}^{F}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\right\|$ (62)
$\leq\sum_{q=1}^{F}\left\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}-{\mathbf{P}}_{N}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\right\|$ (63)
The difference can be further decomposed as
$\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}-{\mathbf{P}}_{N}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\|$
$\leq\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}-{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f_{l-1}^{q}+{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f_{l-1}^{q}-{\mathbf{P}}_{N}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\|$ (64)
$\leq\left\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{n,l-1}^{q}-{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f_{l-1}^{q}\right\|+\left\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f_{l-1}^{q}-{\mathbf{P}}_{N}{\mathbf{h}}_{l}^{pq}({\mathcal{L}}_{\rho})f_{l-1}^{q}\right\|$ (65)
The second term can be bounded with (58) and we denote the bound as $\Delta_{N}$ for simplicity. The first term can be decomposed by Cauchy-Schwartz inequality and non-amplifying of the filter functions as
$\left\|{\mathbf{x}}_{n,l}^{p}-{\mathbf{P}}_{N}f_{l}^{p}\right\|\leq\sum_{q=1}^{F}\Delta_{N}\|{\mathbf{x}}_{n,l-1}^{q}\|+\sum_{q=1}^{F}\|{\mathbf{x}}_{l-1}^{q}-{\mathbf{P}}_{N}f_{l-1}^{q}\|.$ (66)
To solve this recursion, we need to compute the bound for $\|{\mathbf{x}}_{l}^{p}\|$. By normalized Lipschitz continuity of $\sigma$ and the fact that $\sigma(0)=0$, we can get
$\|{\mathbf{x}}_{l}^{p}\|\leq\left\|\sum_{q=1}^{F}{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N}){\mathbf{x}}_{l-1}^{q}\right\|\leq\sum_{q=1}^{F}\left\|{\mathbf{h}}_{l}^{pq}({\mathbf{L}}_{N})\right\|\|{\mathbf{x}}_{l-1}^{q}\|\leq\sum_{q=1}^{F}\|{\mathbf{x}}_{l-1}^{q}\|\leq F^{l-1}\|{\mathbf{x}}\|.$ (67)
Insert this conclusion back to solve the recursion, we can get
$\left\|{\mathbf{x}}_{n,l}^{p}-{\mathbf{P}}_{N}f_{l}^{p}\right\|\leq lF^{l-1}\Delta_{N}\|{\mathbf{x}}\|.$ (68)
Replace $l$ with $L$ we can obtain
$\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{P}}_{N}f)-{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f))\|\leq LF^{L-1}\Delta_{N},$ (69)
when the input graph signal is normalized. By replacing $f={\mathbf{I}}_{N}{\mathbf{x}}$, we can conclude the proof. ∎
Appendix C Local Lipschitz continuity of MNNs
We propose that the outputs of MNN defined in (7) are locally Lipschitz continuous within a certain area, which is stated explicitly as follows.
Proposition 3.
(Local Lipschitz continuity of MNNs) Let MNN be $L$ layers with $F$ features in each layer, suppose the manifold filters are nonamplifying with $|\hat{h}(\lambda)|\leq 1$ and the nonlinearities normalized Lipschitz continuous, then there exists a constant $C$ such that
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\text{dist}(x-y),\quad\text{for all }x,y\in B_{r}({\mathcal{M}}),$ (70)
where $B_{r}({\mathcal{M}})$ is a ball with radius $r$ over ${\mathcal{M}}$.
Proof.
The output of MNN can be written explicitly as
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|=\left|\sigma\left(\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)\right)-\sigma\left(\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right)\right|$ (71)
$\leq\left|\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)-\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right|\leq F\max_{q=1,\cdots,F}\left|{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)-{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right|.$ (72)
We have $f_{L-1}^{q}(x)=\sigma\left(\sum_{p=1}^{F}{\mathbf{h}}_{L-1}^{p}f_{L-2}^{p}(x)\right)$. The process can be repeated recursively, and finally, we can have
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}|{\mathbf{h}}_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f(x)-{\mathbf{h}}_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f(y)|.$ (73)
With $f$ as a $\lambda$-bandlimited manifold signal, we suppose $g=bbh_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f$. As $\langle f,\bm{\phi}_{i}\rangle=0$ for all $i>M$, $g$ is also bandlimited and possesses $M$ spectral components. The gradient can be bounded according to [51] combined with the non-amplifying property of the filter function as
$\|\nabla g\|_{\infty}\leq C\sum_{\lambda_{i}\leq\lambda}\left|\hat{h}(\lambda_{i})\right|^{L}\lambda_{i}^{\frac{d+1}{2}}\|f\|_{\mathcal{M}}\leq C\sum_{\lambda_{i}\leq\lambda}\lambda_{i}^{\frac{d+1}{2}}\|f\|_{\mathcal{M}}$ (74)
From Theorem 4.5 in [52], $g$ is locally Lipschitz continuous as
$|g(x)-g(y)|\leq C^{\prime}\text{dist}(x-y),\quad\text{with }x,y\in B_{r}({\mathcal{M}}),$ (75)
where $B_{r}({\mathcal{M}})$ is a closed ball with radius $r$.
Combining the above, we have the continuity of the output of MNN as
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\text{dist}(x-y),\quad\text{with}x,y\in B_{r}({\mathcal{M}}),$ (76)
which concludes the proof. ∎
Appendix D Proof of Theorem 1
Proof.
To analyze the difference between the empirical risk and statistical risk, we introduce an intermediate term which is the induced version of the sampled MNN output. We define ${\mathbf{I}}_{N}$ as the inducing operator based on the Voronoi decomposition $\{V_{i}\}_{i=1}^{N}$ defined in Section A. This intermediate term is written explicitly as
$\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)={\mathbf{I}}_{N}{\mathbf{P}}_{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)=\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})\mathbbm{1}_{x\in V_{i}},\;\text{for all}\;x\in{\mathcal{M}},$ (77)
where $x_{i}\in X_{N}$ are sampled points from the manifold.
Suppose ${\mathbf{H}}\in\arg\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})$, we have
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})|$ (78)
The difference between $R_{\mathbf{G}}({\mathbf{H}})$ and $R_{\mathcal{M}}({\mathbf{H}})$ can be decomposed as
$\left|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})\right|$
$=\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (79)
$=\Bigg{|}\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)$
$\qquad\qquad+\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\Bigg{|}$ (80)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$
$\qquad\qquad+\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (81)
We analyze the two terms in (81) separately, with the first term bounded based on the convergence of GNN to MNN and the second term bounded with the smoothness of manifold functions.
The first term in (81) can be written as
$\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (82)
$=\frac{1}{N}\left|\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\sum_{i=1}^{N}\ell(\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (83)
$\leq\frac{1}{N}\sum_{i=1}^{N}\left|\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\ell(\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (84)
$\leq\frac{1}{N}\sum_{i=1}^{N}\Big{|}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i}-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})\Big{|}$ (85)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{1}$ (86)
$\leq\frac{1}{\sqrt{N}}LF^{L-1}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (87)
From (82) to (83), we use the definition of induced manifold signal defined in (77). We utilize the Lipschitz continuity assumption on loss function from (84) to (85). From (85) to (86), it depends on the fact that ${\mathbf{x}}$ is a single-entry vector and that $[{\mathbf{y}}]_{i}$ is the value sampled from target manifold function $g$ evaluated on $x_{i}$. Finally the bound depends on the convergence of GNN on the sampled graph to the MNN as stated in Proposition 1.
The second term is decomposed as
$\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (88)
$\leq\left|\sum_{i=1}^{N}\int_{V_{i}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\sum_{i=1}^{N}\int_{V_{i}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (89)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)-\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\right|\text{d}\mu(x)$ (90)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$ (91)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$ (92)
From (88) to (89), it relies on the decomposition of the MNN output over $\{V_{i}\}_{i=1}^{N}$. From (90) to (91), we use the Lipschitz continuity of loss function. From (91) to (92), we use the definition of $\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)$. Proposition 3 indicates that the MNN outputs are Lipschitz continuous within a certain range, which leads to
$\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$
$\leq\sum_{i=1}^{N}\int_{V_{i}}F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}\text{d}\mu(x)$ (93)
$=F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}\sum_{i=1}^{N}\int_{V_{i}}\text{d}\mu(x)$ (94)
$\leq F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}.$ (95)
Combining (87) and (D), we can conclude the proof. ∎
Appendix E Proof of Theorem 2
Proof.
We can write the difference as
$\left|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})\right|$
$\leq\sum_{k=1}^{K}\left|\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)\right|$ (96)
Based on the property of absolute value inequality and the Lipschitz continuity assumption of loss function (Assumption 2), we have
$\left|\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)\right|$
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (97)
We insert an intermediate term $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})$ as the value evaluated on the sampled point $x_{i}$, which leads to
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (98)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|+$
$\qquad\qquad\qquad\qquad\qquad\qquad\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (99)
The first term in (E) can be bounded similarly as (86), which is explicitly written as
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|$ (100)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f_{k})\|_{1}$ (101)
$\leq\frac{1}{\sqrt{N}}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f_{k})\|_{2}$ (102)
$\leq\frac{1}{\sqrt{N}}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (103)
The second term is
$\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (104)
$=\left|\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\text{d}\mu_{k}(x)-\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x)\text{d}\mu_{k}(x)\right|$ (105)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x)\right|\text{d}\mu_{k}(x)$ (106)
$\leq F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}$ (107)
This depends on the Lipschitz continuity of the output manifold function in Proposition 3. ∎
Appendix F Experiment details and further experiments
All experiments were done using a NVIDIA GeForce RTX 3090, and each set of experiments took at most $10$ hours to complete. In total, we run $10$ datasets, which amounts for around $100$ hours of GPU use. All datasets used in this paper are public, and free to use. They can be downloaded using the $pytorch$ package (https://pytorch-geometric.readthedocs.io/en/latest/modules/datasets.html), the ogb package (https://ogb.stanford.edu/docs/nodeprop/) and the Princeton ModelNet project (https://modelnet.cs.princeton.edu/). In total, the datasets occupy around $5$ gb. However, they do not need to be all stored at the same time, as the experiments that we run can be done in series.
F.1 ModelNet10 and ModelNet40 graph classificationtasks
ModelNet10 dataset [43] includes 3,991 meshed CAD models from 10 categories for training and 908 models for testing as Figure 7 shows. ModelNet40 dataset includes 38,400 training and 9,600 testing models as Figure 8 shows. In each model, $N$ points are uniformly randomly selected to construct graphs to approximate the underlying model, such as chairs, tables.
Figure 7: Point cloud models in ModelNet10 with $N=300$ sampled points in each model, corresponding to bathtub, chair, desk, table, toiler, and bed.
Figure 8: Point cloud models from ModelNet40 with $N=300$ sampled points in each model, corresponding to airplane, person, car, guitar, plant, and bottle.
(a) Differences of the outputs of 3-layer GNNs.
(b) Differences of the outputs of 4-layer GNNs.
Figure 9: Graph outputs differences of GNNs with different architectures on ModelNet40 dataset.
The weight function of the constructed graph is determined as (8) with $\epsilon=0.1$. We calculate the Laplacian matrix for each graph as the input graph shift operator. In this experiment, we implement GNNs with different numbers of layers and hidden units with $K=5$ filters in each layer. All the GNN architectures are trained by minimizing the cross-entropy loss. We implement an ADAM optimizer with the learning rate set as 0.005 along with the forgetting factors 0.9 and 0.999. We carry out the training for 40 epochs with the size of batches set as 10. We run 5 random dataset partitions and show the average performances and the standard deviation across these partitions.
F.2 Node classification training details and datasets
In this section, we present the results for node classification. In this paragraph we present the common details for all datasets, we will next delve into each specific detail inside the dataset subsection that follows.
Name Nodes Edges Features
Number
of Classes
Reference
Arxiv $169,343$ $1,166,243$ $128$ $40$ [20, 44]
Cora $2,708$ $10,556$ $1,433$ $7$ [45]
CiteSeer $3,327$ $9,104$ $3,703$ $6$ [45]
PubMed $19,717$ $88,648$ $500$ $3$ [45]
Coauthor Physics $18,333$ $163,788$ $6,805$ $15$ [46]
Coauthor CS $34,493$ $495,924$ $8,415$ $5$ [46]
Amazon-ratings $24,492$ $93,050$ $300$ $5$ [47]
Roman-empire $22,662$ $32,927$ $300$ $18$ [47]
Table 2: Details of the datasets considered in the experiments.
In all datasets, we used the graph convolutional layer GCN, and trained for $1000$ epochs. For the optimizer, we used AdamW, with using a learning rate of $0.01$, and 00 weight decay. We trained using the graph convolutional layer, with a varying number of layers and hidden units. For dropout, we used $0.5$. We trained using the cross-entropy loss. In all cases, we trained $2$ and $3$ layered GNNs.
To compute the linear approximation in the plots, we used the mean squared error estimator of the form
$\mathbf{y}=s*\log(\mathbf{n})+p.$ (108)
Where $s$ is the slope, $p$ is the point, and $\mathbf{n}$ is the vector with the nodes in the training set for each experiment. Note that we repeated each experiment for $10$ independent runs. In all experiments, we compute the value of $s$ and $p$ that minimize the mean square error over the mean of the experiment runs, and we compute the Pearson correlation index over those values.
Our experiment shows that our bound shows the same rate dependency as the experiments. That is to say, in the logarithmic scale, the generalization gap of GNNs is linear with respect to the logarithm of the number of nodes. In most cases, the Pearson correlation index is above $0.9$ in absolute value, which indicates a strong linear relationship. We noticed that the linear relationship changes the slope in the overfitting regime, and in the non-overfitting regime. That is to say, when the GNN is overfitting the training set, the generalization gap decreases at a much slower rate than it does with the GNN does not have the capacity to do so. Therefore, in the case in which the GNN overfits the training set for all nodes when computed $s$ using all the samples in the experiment. On the other hand, when the number of nodes is large enough that the GNN cannot overfit the training set, then we computed the $s$ and $p$ with the nodes in the non overfitting regime.
F.2.1 Arxiv dataset
For this datasets, we trained $2,3,4$ layered GNN. We also used a learning rate scheduler ReduceLROnPlateau with mode min, factor $0.5$, patience $100$ and a minimum learning rate of $0.001$.
(a) Two Layers
(b) Three Layers
(c) Four Layers
Figure 10: Generalization gap for the OGBN-Arxiv dataset on the accuracy as a function of the number of nodes in the training set.
(a) Two Layers
(b) Three Layers
(c) Four Layers
Figure 11: Generalization gap for the OGBN-arxiv dataset on the loss (cross-entropy) as a function of the number of nodes in the training set.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $64$ $-6.301e-01$ $003.621e+$ 00 $-9.980e-01$
Accuracy $2$ $128$ $-6.034e-01$ $003.663e+$ 00 $-9.985e-01$
Accuracy $2$ $256$ $-5.347e-01$ $003.493e+$ 00 $-9.952e-01$
Accuracy $2$ $512$ $-5.328e-01$ $003.605e+$ 00 $-9.975e-01$
Accuracy $3$ $64$ $-6.271e-01$ $003.600e+$ 00 $-9.987e-01$
Accuracy $3$ $128$ $-5.730e-01$ $003.567e+$ 00 $-9.970e-01$
Accuracy $3$ $256$ $-4.986e-01$ $003.393e+$ 00 $-9.910e-01$
Accuracy $3$ $512$ $-4.529e-01$ $003.315e+$ 00 $-9.934e-01$
Accuracy $4$ $64$ $-5.343e-01$ $003.236e+$ 00 $-9.971e-01$
Accuracy $4$ $128$ $-5.096e-01$ $003.299e+$ 00 $-9.987e-01$
Accuracy $4$ $256$ $-4.827e-01$ $003.337e+$ 00 $-9.920e-01$
Accuracy $4$ $512$ $-4.264e-01$ $003.229e+$ 00 $-9.927e-01$
Loss $2$ $64$ $-6.853e-01$ $002.265e+$ 00 $-9.975e-01$
Loss $2$ $128$ $-6.562e-01$ $002.311e+$ 00 $-9.988e-01$
Loss $2$ $256$ $-5.907e-01$ $002.174e+$ 00 $-9.968e-01$
Loss $2$ $512$ $-5.848e-01$ $002.280e+$ 00 $-9.989e-01$
Loss $3$ $64$ $-6.739e-01$ $002.228e+$ 00 $-9.980e-01$
Loss $3$ $128$ $-6.229e-01$ $002.224e+$ 00 $-9.976e-01$
Loss $3$ $256$ $-5.581e-01$ $002.111e+$ 00 $-9.942e-01$
Loss $3$ $512$ $-5.141e-01$ $002.057e+$ 00 $-9.955e-01$
Loss $4$ $64$ $-6.039e-01$ $001.964e+$ 00 $-9.980e-01$
Loss $4$ $128$ $-5.701e-01$ $002.014e+$ 00 $-9.991e-01$
Loss $4$ $256$ $-5.379e-01$ $002.051e+$ 00 $-9.951e-01$
Loss $4$ $512$ $-4.810e-01$ $001.957e+$ 00 $-9.937e-01$
Table 3: Details of the linear approximation of the Arxiv Dataset. Note that in this case, we used only the values of the generalization gap whose training error is below $95\%$.
(a) Accuracy
(b) Loss
Figure 12: Values of slope (a) and point (b) corresponding to the linear fit ($a*\texttt{log}(N)+b$) of Figures 11 and 10.
F.2.2 Cora dataset
For the Cora dataset, we used the standard one, which can be obtained running torch_geometric.datasets.Planetoid(root="./data",name=’Cora’).
(a) Generalization Gap.
(b) $2$ Layers
(c) $3$ Layers
Figure 13: Generalization gap, testing, and training losses with respect to the number of nodes in the Cora dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $16$ $-2.839e-01$ $002.022e+$ 00 $-9.803e-01$
Accuracy $2$ $32$ $-2.917e-01$ $002.014e+$ 00 $-9.690e-01$
Accuracy $2$ $64$ $-3.006e-01$ $002.021e+$ 00 $-9.686e-01$
Accuracy $3$ $16$ $-2.656e-01$ $001.996e+$ 00 $-9.891e-01$
Accuracy $3$ $32$ $-2.637e-01$ $002.008e+$ 00 $-9.679e-01$
Accuracy $3$ $64$ $-2.581e-01$ $001.981e+$ 00 $-9.870e-01$
Loss $2$ $16$ $-3.631e-01$ $019.406e-$ 01 $-9.250e-01$
Loss $2$ $32$ $-4.228e-01$ $019.638e-$ 01 $-9.657e-01$
Loss $2$ $64$ $-4.991e-01$ $001.067e+$ 00 $-9.776e-01$
Loss $3$ $16$ $-4.131e-01$ $001.276e+$ 00 $-9.753e-01$
Loss $3$ $32$ $-4.605e-01$ $001.385e+$ 00 $-9.730e-01$
Loss $3$ $64$ $-4.589e-01$ $001.455e+$ 00 $-9.756e-01$
Table 4: Details of the linear approximation of the Cora Dataset. Note that in this case we used all the values given that the training accuracy is $100\%$ for all nodes.
F.2.3 CiteSeer dataset
For the CiteSeer dataset, we used the standard one, which can be obtained running torch_geometric.datasets.Planetoid(root="./data",name=’CiteSeer’).
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
Figure 14: Generalization gap, testing, and training losses with respect to the number of nodes in the CiteSeer dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $16$ $-1.699e-01$ $001.972e+$ 00 $-9.518e-01$
Accuracy $2$ $32$ $-1.856e-01$ $001.978e+$ 00 $-9.714e-01$
Accuracy $2$ $64$ $-1.749e-01$ $001.966e+$ 00 $-9.534e-01$
Accuracy $3$ $16$ $-1.585e-01$ $001.956e+$ 00 $-9.721e-01$
Accuracy $3$ $32$ $-1.659e-01$ $001.963e+$ 00 $-9.721e-01$
Accuracy $3$ $64$ $-1.658e-01$ $001.967e+$ 00 $-9.702e-01$
Loss $2$ $16$ $-1.049e-01$ $017.757e-$ 01 $-5.924e-01$
Loss $2$ $32$ $-1.762e-01$ $017.646e-$ 01 $-7.981e-01$
Loss $2$ $64$ $-2.186e-01$ $018.384e-$ 01 $-9.120e-01$
Loss $3$ $16$ $-1.802e-01$ $001.169e+$ 00 $-8.345e-01$
Loss $3$ $32$ $-1.629e-01$ $001.200e+$ 00 $-8.767e-01$
Loss $3$ $64$ $-5.917e-02$ $001.283e+$ 00 $-2.562e-01$
Table 5: Details of the linear approximation of the CiteSeer Dataset. Note that in this case we used all the values given that the training accuracy is $100\%$ for all nodes.
F.2.4 PubMed dataset
For the PubMed dataset, we used the standard one, which can be obtained running torch_geometric.datasets.Planetoid(root="./data",name=’PubMed’).
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
Figure 15: Generalization gap, testing, and training losses with respect to the number of nodes in the PubMed dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $16$ $-2.523e-01$ $001.834e+$ 00 $-9.942e-01$
Accuracy $2$ $32$ $-2.433e-01$ $001.812e+$ 00 $-9.583e-01$
Accuracy $2$ $64$ $-2.764e-01$ $001.869e+$ 00 $-9.761e-01$
Accuracy $3$ $16$ $-2.748e-01$ $001.844e+$ 00 $-9.910e-01$
Accuracy $3$ $32$ $-2.661e-01$ $001.861e+$ 00 $-9.712e-01$
Accuracy $3$ $64$ $-2.558e-01$ $001.827e+$ 00 $-9.890e-01$
Loss $2$ $16$ $-4.166e-01$ $017.695e-$ 01 $-9.718e-01$
Loss $2$ $32$ $-4.733e-01$ $017.852e-$ 01 $-9.137e-01$
Loss $2$ $64$ $-4.368e-01$ $017.547e-$ 01 $-9.718e-01$
Loss $3$ $16$ $-4.424e-01$ $001.067e+$ 00 $-9.549e-01$
Loss $3$ $32$ $-5.518e-01$ $001.223e+$ 00 $-9.655e-01$
Loss $3$ $64$ $-5.246e-01$ $001.169e+$ 00 $-9.632e-01$
Table 6: Details of the linear approximation of the PubMed Dataset. Note that in this case we used all the values given that the training accuracy is $100\%$ for all nodes.
F.2.5 Coauthors CS dataset
For the CS dataset, we used the standard one, which can be obtained running torch_geometric.datasets.Coauthor(root="./data", name=’CS’). In this case, given that there are no training and testing sets, we randomly partitioned the datasets and used $90\%$ of the samples for training and the remaining $10\%$ for testing.
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
Figure 16: Generalization gap, testing, and training losses with respect to the number of nodes in the CS dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
(a) Linear fit for accuracy generalization gap
(b) Linear fit for loss generalization gap
Figure 17: Generalization gaps as a function of the number of nodes in the training set in the CS dataset.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $32$ $-2.138e-01$ $001.659e+$ 00 $-9.007e-01$
Accuracy $2$ $64$ $-2.250e-01$ $001.685e+$ 00 $-8.969e-01$
Accuracy $3$ $32$ $-1.979e-01$ $001.695e+$ 00 $-9.009e-01$
Accuracy $3$ $64$ $-1.862e-01$ $001.646e+$ 00 $-8.980e-01$
Loss $2$ $32$ $-2.523e-01$ $016.273e-$ 01 $-8.244e-01$
Loss $2$ $64$ $-2.933e-01$ $017.762e-$ 01 $-7.925e-01$
Loss $3$ $32$ $-3.558e-01$ $001.207e+$ 00 $-8.924e-01$
Loss $3$ $64$ $-3.560e-01$ $001.256e+$ 00 $-8.568e-01$
Table 7: Details of the linear approximation of the CS Dataset. Note that in this case we used all the values given that the training accuracy is $100\%$ for all nodes.
F.2.6 Coauthors Physics dataset
For the Physics dataset, we used the standard one, which can be obtained running torch_geometric.datasets.Coauthor(root="./data", name=’Physics’). In this case, given that there are no training and testing sets, we randomly partitioned the datasets and used $90\%$ of the samples for training and the remaining $10\%$ for testing.
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
Figure 18: Generalization gap, testing, and training losses with respect to the number of nodes in the Physics dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
(a) Linear fit for accuracy generalization gap
(b) Linear fit for loss generalization gap
Figure 19: Generalization Gaps as a function of the number of nodes in the training set in the Physics dataset.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $32$ $-1.524e-01$ $001.235e+$ 00 $-9.064e-01$
Accuracy $2$ $64$ $-1.478e-01$ $001.218e+$ 00 $-9.145e-01$
Accuracy $3$ $32$ $-1.227e-01$ $001.190e+$ 00 $-9.328e-01$
Accuracy $3$ $64$ $-1.268e-01$ $001.200e+$ 00 $-8.826e-01$
Loss $2$ $32$ $-1.111e-01$ $-5.257e-02$ $-7.591e-01$
Loss $2$ $64$ $-9.684e-02$ $-7.335e-02$ $-7.696e-01$
Loss $3$ $32$ $-1.410e-01$ $012.875e-$ 01 $-8.280e-01$
Loss $3$ $64$ $-1.068e-01$ $012.388e-$ 01 $-7.679e-01$
Table 8: Details of the linear approximation of the Physics Dataset. Note that in this case we used all the values given that the training accuracy is $100\%$ for all nodes.
F.2.7 Heterophilous Amazon ratings dataset
For the Amazon dataset, we used the standard one, which can be obtained running torch_geometric.datasets.HeterophilousGraphDataset(root="./data", name=’Amazon’). In this case, we used the $10$ different splits that the dataset has assigned.
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
Figure 20: Generalization gap, testing, and training losses with respect to the number of nodes in the Amazon dataset. The top row is in accuracy, and the bottom row is the cross-entropy loss.
(a) Linear fit for accuracy generalization gap
(b) Linear fit for loss generalization gap
Figure 21: Generalization Gaps as a function of the number of nodes in the training set in the Amazon dataset.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $32$ $-7.693e-01$ $004.236e+$ 00 $-9.914e-01$
Accuracy $2$ $64$ $-7.788e-01$ $004.404e+$ 00 $-9.972e-01$
Accuracy $3$ $32$ $-7.268e-01$ $004.101e+$ 00 $-9.868e-01$
Accuracy $3$ $64$ $-7.354e-01$ $004.257e+$ 00 $-9.921e-01$
Loss $2$ $32$ $-1.086e+00$ $003.971e+$ 00 $-9.968e-01$
Loss $2$ $64$ $-1.096e+00$ $004.189e+$ 00 $-9.985e-01$
Loss $3$ $32$ $-1.134e+00$ $004.339e+$ 00 $-9.965e-01$
Loss $3$ $64$ $-1.154e+00$ $004.629e+$ 00 $-9.991e-01$
Table 9: Details of the linear approximation of the Amazon Dataset. Note that in this case we used only the values of the generalization gap whose training error is below $95\%$.
F.2.8 Heterophilous Roman Empire dataset
For the Roman dataset, we used the standard one, which can be obtained running torch_geometric.datasets.HeterophilousGraphDataset(root="./data", name=’Roman’). In this case, we used the $10$ different splits that the dataset has assigned.
(a) Generalization Gap
(b) $2$ Layers
(c) $3$ Layers
(a) Linear fit for accuracy generalization gap
(b) Linear fit for loss generalization gap
Figure 23: Generalization Gaps as a function of the number of nodes in the training set in the Roman dataset.
Type Lay. Feat. Slope Point
Pearson
Correlation
Coefficient
Accuracy $2$ $32$ $-8.408e-01$ $004.644e+$ 00 $-9.963e-01$
Accuracy $2$ $64$ $-7.435e-01$ $004.477e+$ 00 $-1.000e+00$
Accuracy $3$ $32$ $-9.476e-01$ $005.049e+$ 00 $-9.956e-01$
Accuracy $3$ $64$ $-9.145e-01$ $005.182e+$ 00 $-1.000e+00$
Loss $2$ $32$ $-1.006e+00$ $003.829e+$ 00 $-9.992e-01$
Loss $2$ $64$ $-9.656e-01$ $003.915e+$ 00 $-1.000e+00$
Loss $3$ $32$ $-1.244e+00$ $004.764e+$ 00 $-9.994e-01$
Loss $3$ $64$ $-1.225e+00$ $005.011e+$ 00 $-1.000e+00$
Table 10: Details of the linear approximation of the Roman Dataset. Note that in this case we used only the values of the generalization gap whose training error is below $95\%$
Appendix G Further references
Graphon theory
Related to the manifold model we consider, some works construct graphs sampled from graphons, which can be seen as a random limit graph model. This line of work has concentrated on their convergence, stability and transferability properties [53, 54, 55]. Graphon is also used as a tool for pooling in GNNs in [56]. Albeit useful as a model, the graphon presents several limitations compared to the manifold model we use. Firstly, the graphon model assumes an infinite degree at every node [57], which is not the case in the manifold model. Secondly, the graphon presents little intuition on the underlying model, outside of the stochastic block model graphon, it is not trivial to visualize a graphon. On the other hand, manifolds are more interpretable, as can be the case on the sphere, or on other shapes (see Figure 2). Lastly, the manifold model is a more realistic assumption in the types of models that it can characterize.
Manifold Theory
We consider an underlying manifold model from which we sample to obtain graphs, which are later the inputs to the GNN. Related to us, some works have used the sampled graphs the infer properties about the manifold itself. In particular, some works have set out to test the manifold assumption of the data [58], the manifold dimension [59] and show the complexity of doing so [60, 61]. In terms of prediction and classification using manifolds, and manifold data, several algorithms and methods have been proposed – notable the Isomap algorithm [62, 63, 64] and other manifold learning techniques [65]. In all, these techniques utilize the manifold assumption to infer properties of the manifold, which differs from our goal of providing a generalization bound for GNNs.
Transferability of GNNs
Transferability of GNNs has been analyzed in many works by comparing the output differences of GNNs on different sizes of graphs when graphs converge to a limit model without statistical generalization analysis. In [15, 53, 54], the transferability of GNNs on graphs generated from graphon models is analyzed with difference bounds of GNN outputs proved. In [66], the authors show how increasing the size of the graph as the GNN learns, generalizes to the large-scale graph. In [67], the authors analyze the transferability of GNNs on graphs generated from a general topological space while the authors in [35] focus on graphs generated from manifolds. In [68], the authors propose a novel graphop operator as a limit model for both dense and sparse graphs with a transferability result proved. In [69] and [70], the authors study the transfer learning of GNNs with measurements of distance between graphs without a limit assumption. In [71], a transferable graph transformer is proposed and proved with empirical results.
Appendix H Low Pass Filter Assumption
In the main results, we assume that the GNN and MNN are low-pass filters. We believe this is a mild assumption. A high frequency on the graph/manifold implies that there exist signals (eigenvectors/functions) that can have large variations in adjacent entries. This naturally generates instabilities, which are more difficult to learn. It is expected certain amount of homogeneity within the neighborhood of a node, translates into low-frequency signals. It has been shown in practice that several real-world physical models lead to low-pass filters such as opinion dynamics [72], econometrics [73], and graph signal processing [74].
Several other learning techniques rely on low-pass filters. Principal Component Analysis (PCA), can be seen as low-pass filters applied to the correlation matrix. Isomap, is another example of low pass filtering, but in this case applied to the Laplacian matrix. In all, we believe that the low pass filter assumption is not a hard assumption, given its empirical verification, and the fact that several other techniques that work in practice utilize it.